OpenStudy (anonymous):

Essay; show all work. Consider that the age, x, of a unicorn in human equivalent years can be given by the formula f(x) = - 0.001518x4 + 0.097326x3 – 1.2367x2 + 10.46x + 2.914. When a unicorn is 2.5 years old, what is its age in human equivalent years? What about when it is 12 years old?

OpenStudy (anonymous):

does anyone understand this ? I need help asap

OpenStudy (amistre64):

substitute 2.5 for x; since x = 2.5 then do the same again but for x = 12

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

f(2.5)=-0.001518*2.5^4 + 0.097326*2.5^3 – 1.2367*2.5^2 + 10.46*2.5 + 2.914=22.8 2.5 unicorn years is equivalent to 23 unicorn years (2 significant figures) f(12)=-0.001518*12^4 + 0.097326*12^3 – 1.2367*12^2 + 10.46*12 + 2.914=87.1 12 unicorn years is equivalent to 87 unicorn years (2 significant figures)

OpenStudy (anonymous):

As amistre said, it is a substitution problem.